OpenStudy (anonymous):

Two lines, A and B, are represented by the equations given below: Line A: x + y = 2 Line B: 2x + y = 4 Which statement is true about the solution to the set of equations? There are infinitely many solutions. There are two solutions. There is one solution. There is no solution.

OpenStudy (anonymous):

@brandonford7 @Sachintha

OpenStudy (anonymous):

@snugbug505

OpenStudy (anonymous):

is it one solution?

OpenStudy (anonymous):

thanks!!! @Sachintha

OpenStudy (anonymous):

Please don't give out direct answers… it's against the CoC. @Sachintha

OpenStudy (anonymous):

Try explaining it.

OpenStudy (sachintha):

Anyway my answer is wrong :D

OpenStudy (anonymous):

so one solution?

OpenStudy (anonymous):

huh?

OpenStudy (sachintha):

How did you get it?

OpenStudy (anonymous):

guessed lol

OpenStudy (sachintha):

Subtract equation A from B so that you can get a value for x.

OpenStudy (anonymous):

-2

OpenStudy (sachintha):

Nope. \[2x+y-(x+y)=4-2\] \[2x+y-x-y=2\] \[x=2\]

OpenStudy (anonymous):

ohhh... so its one solution?

OpenStudy (sachintha):

For y you just substitute the value for x.

OpenStudy (sachintha):

Yeah :)

OpenStudy (anonymous):

thanks so much